banani:

Mrs. Williams is deciding between two field trips for her class. The Science Center charges $125 plus $3 per student. The Dino Discovery Museum simply charges $8 per student. For how many students will the Science Center charge less than the Dino Discovery Museum?

surjithayer:

let x be the number of students when the charges are equal. 8x=125+3x 8x-3x=125 5x=125 x=125/5=25 if x=24 charges for science center=125+3*24=125+72=197 $ charges for Dino Discovery Museum=24*8=$192 if x=26 charges for science center=125+26*3=125+78=$203 charges for Dino Discovery Museum=26*8=$208 so if number of students >25 then charges for science center is less than the charges for Dino Discovery Museum.